OpenStudy (mackenzie2013):

A square has an area of 16 square millimeters. What is the length of each side of the square? 2 mm 4 mm 8 mm 12 mm Please explain the steps! :)

OpenStudy (ikram002p):

square sides are the same 

OpenStudy (ikram002p):

area of the square is \( (side)^2\)

OpenStudy (ikram002p):

for this side area=\(x^2\)= 16 

OpenStudy (ikram002p):

so \(x^2=16\) then \(x=\sqrt 16=4\)

OpenStudy (anonymous):

since a square has equal lengths and the area of a square is length time width. area=16 root16=4